Consider the GDP data in the table below for some economy for the years 2012 through 2015. Real GDP is calculated using the chain-weighted method. 2012 1,250 1,000 2013 2015 2,800 2,030 2014 Nominal GDP Real GDP Suppose you were told that the value of A is 1,600, the value of B is 2,240, the value of C is 1,450, and the value of D is 1,624. If you were to use the year 2015 as the base 3? (Do not use commas or decimals for your answer) year,what is the value of chain-weighted real GDP for 201
Answer #1

Base

2012

2013

2014

2015

Nominal GDP

1250

1600

2240

2800

Real GDP

1000

1450

1624

2030

GDP Deflator

125.00

110.34

137.93

137.93

Using 2015 as the base year, we know that Real GDP is equal to nominal GDP. Thus Real GDP in 2015 is $2030. This gives us the starting point for the chain-weighted method of calculating real GDP.

To calculate chain-weighted Real GDP for 2013 we need the following four pieces of information:
2015 Real GDP at 2015 prices: $2030
2013 Real GDP at 2013 prices: $1450.
2015 Real GDP at 2013 prices =.2030/(1.1) = 1845.45
2013 Real GDP at 2015 prices = 1450/(1.38) = 1050.72
Calculate the growth rate of GDP with 2015 prices:
((2030-1050.72)/2030)*100 = 48.24%,
Then the growth rate of GDP using 2013 prices:
((1845.45-1450)/1450)*100 = 27.27%.
The next step is to average the two growth rates: (48.24+27.27)/2 = 37.75%.
So to calculate 2013 Real GDP we multiply 2015 real GDP by this growth rate:
(2030+(2030*37.75%)) = $2796